# Grassroots Bonds: A Grassroots Foundation for Market Liquidity

### 14 Mar 2026 v1

@@ Appendix B Bond Agent Source Code
        TradeResp, OfferedBonds?,
        Selected?, Remaining?,
        UserIn?, NetIn?, Outs?, NextSerial?).
trade_dispatch(normal, Id, From, WantSpec, OfferedBonds,
    TradeResp?,
    Holdings, UserIn, NetIn, Outs, NextSerial) :-
@@
    agent(Id?, UserIn?, NetIn?, Outs1?, Holdings?,
        NextSerial?).
%% REDEMPTION_RESULT
procedure redemption_result(Constant?,
    Constant?, Constant?, TradeResponse,
    BondList?, BondList?, BondList?,
    UserInStream?, NetInStream?,
    OutputsList?, Constant?).
%% Can fill -- accept with requested bonds
@@
    trade_accept(Selected?), OfferedBonds,
    Selected, Remaining,
    UserIn, NetIn, Outs, NextSerial) :-
@@
        msg(agent, ’_user’,
            trade_completed(From?)),
@@
    OfferedBonds,
    Selected, Remaining,
@@
    redemption_reject(Id?, From?,
        OrigHoldings?, Menu,
        UserIn?, NetIn?, Outs?, NextSerial?).
procedure redemption_reject(Constant?,
    Constant?, BondList?, BondList,
    UserInStream?, NetInStream?,
    OutputsList?, Constant?).
@@
        msg(agent, ’_user’,
            trade_failed(From?)),
@@
%% ESCROW -- Concurrent race between timer and cancel signal
procedure escrow(Constant?, BondList?, EscrowCancel?,
    EscrowBenResult, EscrowDepResult).
escrow(T, Bonds, _,
    escrow_bonds(Bonds?), escrow_expired) :-
    wait_until(T?) | true.
escrow(_, Bonds, cancel,
    escrow_cancelled, escrow_bonds(Bonds?)).
%% INJECT_ESCROW_BEN_RESULT
procedure inject_escrow_ben_result(
    EscrowBenResult?, Constant?,
    UserInStream?, UserInStream).
inject_escrow_ben_result(escrow_bonds(Bonds), From, Ys,
    [escrow_ben_released(From?, Bonds?)|Ys?]) :-
    ground(From?), ground(Bonds?) | true.
inject_escrow_ben_result(escrow_cancelled, From, Ys,
    [escrow_ben_cancelled(From?)|Ys?]) :-
    ground(From?) | true.
inject_escrow_ben_result(Resp, From, [Y|Ys], [Y?|Ys1?]) :-
    inject_escrow_ben_result(Resp?, From?, Ys?, Ys1).
%% INJECT_ESCROW_DEP_RESULT
procedure inject_escrow_dep_result(
    EscrowDepResult?, Constant?,
    UserInStream?, UserInStream).
inject_escrow_dep_result(escrow_bonds(Bonds), Target, Ys,
    [escrow_dep_returned(Target?, Bonds?)|Ys?]) :-
    ground(Target?), ground(Bonds?) | true.
inject_escrow_dep_result(escrow_expired, Target, Ys,
    [escrow_dep_expired(Target?)|Ys?]) :-
    ground(Target?) | true.
inject_escrow_dep_result(Resp, Target, [Y|Ys], [Y?|Ys1?]) :-
    inject_escrow_dep_result(Resp?, Target?, Ys?, Ys1).
%% DO_DEPOSIT_ESCROW
procedure do_deposit_escrow(Constant?, Constant?, LotList?,
    Constant?, BondList?, UserInStream?, NetInStream?,
    OutputsList?, Constant?).
do_deposit_escrow(Id, Target, GiveSpec,
    ReleaseTime, Holdings,
    UserIn, NetIn, Outs, NextSerial) :-
@@
procedure do_deposit_escrow_result(Constant?, Constant?,
    Constant?, Constant?, BondList?, BondList?,
    UserInStream?, NetInStream?, OutputsList?, Constant?).
@@
    Selected, Remaining, UserIn, NetIn, Outs, NextSerial) :-
@@
    agent(Id?, UserIn1?, NetIn?, Outs2?, Remaining?,
        NextSerial?).
do_deposit_escrow_result(fail, Id, Target, _, Selected,
    Remaining, UserIn, NetIn, Outs, NextSerial) :-
@@
    agent(Id?, UserIn?, NetIn?, Outs1?, OrigHoldings?,
        NextSerial?).
%% BIND_ESCROW_CANCEL
procedure bind_escrow_cancel(EscrowCancel).
bind_escrow_cancel(cancel).
%% AGENT/6 -- Bond agent
procedure agent(Constant?, UserInStream?, NetInStream?,
    OutputsList?, BondList?, Constant?).
%% --- Phase 1: Mint / Balance / Done ---
agent(Id, [msg(’_user’, Id1, mint(K, Maturity))|UserIn],
@@
    create_bonds(Id?, Maturity?, K?, NextSerial?, NewBonds),
@@
    agent(Id?, UserIn?, NetIn?, Outs1?, NewHoldings?,
        NewNextSerial?).
@@
        msg(agent, ’_user’, balance_report(Holdings?)),
@@
    agent(Id?, UserIn?, NetIn?, Outs1?, Holdings?,
        NextSerial?).
agent(Id, [msg(’_user’, Id1, done)|_], _, Outs, _, _) :-
    Id? =?= Id1? |
    close_outputs(Outs?).
%% --- Phase 2: Social graph (cold-call befriending) ---
agent(Id, [msg(’_user’, Id1, connect(Target))|UserIn],
@@
    lookup_send(’_net’, msg(Target?, intro(Id?, Resp)),
@@
    inject_msg(Resp?, Target?, Id?, UserIn?, UserIn1),
    agent(Id?, UserIn1?, NetIn?, Outs1?, Holdings?,
        NextSerial?).
agent(Id,
    [msg(’_user’, Id1,
@@
    agent(Id?, UserIn?, NetIn1?, Outs1?, Holdings?,
        NextSerial?).
agent(Id, [msg(From, Id1, response(Resp))|UserIn],
@@
    agent(Id?, UserIn?, NetIn1?, Outs1?, Holdings?,
        NextSerial?).
agent(Id, UserIn, [msg(Id1, intro(From, Resp?))|NetIn],
@@
    agent(Id?, UserIn?, NetIn?, Outs1?, Holdings?,
        NextSerial?).
%% --- Phase 7: Trade ---
agent(Id,
    [msg(’_user’, Id1,
@@
        Holdings?, UserIn?, NetIn?, Outs?, NextSerial?).
agent(Id, [trade_complete(From, TheirBonds)|UserIn],
@@
    ground(Id?), ground(From?), ground(TheirBonds?) |
@@
    agent(Id?, UserIn?, NetIn?, Outs1?, NewHoldings?,
        NextSerial?).
agent(Id, [trade_returned_bonds(From, OurBonds)|UserIn],
@@
    ground(Id?), ground(From?), ground(OurBonds?) |
@@
    agent(Id?, UserIn?, NetIn?, Outs1?, NewHoldings?,
        NextSerial?).
agent(Id,
    [trade_returned_bonds_menu(
        From, OurBonds, Menu)|UserIn],
@@
    agent(Id?, UserIn?, NetIn?, Outs1?, NewHoldings?,
        NextSerial?).
agent(Id, UserIn,
    [msg(From, Id1,
@@
    classify_trade(Id?, WantSpec?, OfferedBonds?,
        TradeClass),
    trade_dispatch(TradeClass?, Id?, From?, WantSpec?,
        OfferedBonds?, TradeResp,
        Holdings?, UserIn?, NetIn?, Outs?, NextSerial?).
agent(Id, [msg(’_user’, Id1, accept_trade(From,
@@
    handle_trade_fill(Status?, Id?, From?, TradeResp,
        OfferedBonds?, Selected?, Remaining?,
        UserIn?, NetIn?, Outs?, NextSerial?).
agent(Id, [msg(’_user’, Id1, reject_trade(From,
    trade_pending(TradeResp?, _, OfferedBonds)))|UserIn],
@@
    Id? =?= Id1?, ground(From?), ground(OfferedBonds?) |
@@
    agent(Id?, UserIn?, NetIn?, Outs?, Holdings?,
        NextSerial?).
%% --- Phase 8: Escrow ---
agent(Id,
    [msg(’_user’, Id1,
@@
        ReleaseTime?, Holdings?,
        UserIn?, NetIn?, Outs?, NextSerial?).
@@
        msg(agent, ’_user’, escrow_expired(Target?)),
@@
    agent(Id?, UserIn?, NetIn?, Outs1?, Holdings?,
        NextSerial?).
@@
    ground(Id?), ground(Target?), ground(OurBonds?) |
@@
        msg(agent, ’_user’, escrow_returned(Target?)),
@@
    agent(Id?, UserIn?, NetIn?, Outs1?, NewHoldings?,
        NextSerial?).
agent(Id, [msg(’_user’, Id1, cancel_escrow(Target,
@@
    agent(Id?, UserIn?, NetIn?, Outs?, Holdings?,
        NextSerial?).
agent(Id, UserIn,
    [msg(From, Id1,
@@
    Id? =?= Id1?, ground(From?), ground(ReleaseTime?) |
@@
    agent(Id?, UserIn1?, NetIn?, Outs1?, Holdings?,
        NextSerial?).
agent(Id, [escrow_ben_released(From, Bonds)|UserIn],
@@
        msg(agent, ’_user’, escrow_released(From?)),
@@
    agent(Id?, UserIn?, NetIn?, Outs1?, NewHoldings?,
        NextSerial?).
@@
        msg(agent, ’_user’, escrow_cancelled(From?)),
@@
    agent(Id?, UserIn?, NetIn?, Outs1?, Holdings?,
        NextSerial?).
%% --- Termination ---
agent(_, [], _, Outs, _, _) :- close_outputs(Outs?).

### 17 Mar 2026 v2

@@ Appendix B Bond Agent Source Code
        TradeResp, OfferedBonds?, Selected?,
        Remaining?, UserIn?, NetIn?, Outs?,
        NextSerial?).
trade_dispatch(normal, Id, From, WantSpec,
    OfferedBonds, TradeResp?, Holdings, UserIn,
    NetIn, Outs, NextSerial) :-
@@
    agent(Id?, UserIn?, NetIn?, Outs1?,
        Holdings?, NextSerial?).
%% Redemption result
procedure redemption_result(Constant?, Constant?,
    Constant?, TradeResponse, Stream(Bond)?,
    Stream(Bond)?, Stream(Bond)?,
    Stream(UserInMsg)?, Stream(NetInMsg)?,
    Stream(OutputEntry)?, Constant?).
@@
    trade_accept(Selected?), OfferedBonds, Selected,
    Remaining, UserIn, NetIn, Outs, NextSerial) :-
@@
        msg(agent, ’_user’, trade_completed(From?)),
@@
    OfferedBonds, Selected, Remaining,
@@
    redemption_reject(Id?, From?, OrigHoldings?,
        Menu, UserIn?, NetIn?, Outs?, NextSerial?).
procedure redemption_reject(Constant?, Constant?,
    Stream(Bond)?, Stream(Bond),
    Stream(UserInMsg)?, Stream(NetInMsg)?,
    Stream(OutputEntry)?, Constant?).
@@
        msg(agent, ’_user’, trade_failed(From?)),
@@
%% Escrow — select bonds and create escrow or fail
procedure do_deposit_escrow(Constant?, Constant?,
    Stream(Lot)?, Constant?, Stream(Bond)?,
    Stream(UserInMsg)?, Stream(NetInMsg)?,
    Stream(OutputEntry)?, Constant?).
do_deposit_escrow(Id, Target, GiveSpec, ReleaseTime,
    Holdings, UserIn, NetIn, Outs, NextSerial) :-
@@
procedure do_deposit_escrow_result(Constant?,
    Constant?, Constant?, Constant?, Stream(Bond)?,
    Stream(Bond)?, Stream(UserInMsg)?,
    Stream(NetInMsg)?, Stream(OutputEntry)?,
    Constant?).
@@
    Selected, Remaining, UserIn, NetIn, Outs,
    NextSerial) :-
@@
    agent(Id?, UserIn1?, NetIn?, Outs2?,
        Remaining?, NextSerial?).
do_deposit_escrow_result(fail, Id, Target, _,
    Selected, Remaining, UserIn, NetIn, Outs,
    NextSerial) :-
@@
    agent(Id?, UserIn?, NetIn?, Outs1?,
        OrigHoldings?, NextSerial?).
%% Bond agent — exported entry point
exported procedure agent(Constant?,
    Stream(UserInMsg)?, Stream(NetInMsg)?,
    Stream(OutputEntry)?, Stream(Bond)?, Constant?).
%% Mint
agent(Id,
    [msg(’_user’, Id1, mint(K, Maturity))|UserIn],
@@
    create_bonds(Id?, Maturity?, K?, NextSerial?,
        NewBonds),
@@
    agent(Id?, UserIn?, NetIn?, Outs1?,
        NewHoldings?, NewNextSerial?).
%% Balance
@@
        msg(agent, ’_user’,
            balance_report(Holdings?)),
@@
    agent(Id?, UserIn?, NetIn?, Outs1?,
        Holdings?, NextSerial?).
%% Done
agent(Id, [msg(’_user’, Id1, done)|_], _, Outs,
    _, _) :-
    Id? =?= Id1? | close_outputs(Outs?).
%% Connect (cold call)
agent(Id,
    [msg(’_user’, Id1, connect(Target))|UserIn],
@@
    lookup_send(’_net’,
        msg(Target?, intro(Id?, Resp)),
@@
    inject_msg(Resp?, Target?, Id?,
        UserIn?, UserIn1),
    agent(Id?, UserIn1?, NetIn?, Outs1?,
        Holdings?, NextSerial?).
%% Decision on cold-call
agent(Id, [msg(’_user’, Id1,
@@
    agent(Id?, UserIn?, NetIn1?, Outs1?,
        Holdings?, NextSerial?).
%% Cold-call response (injected)
agent(Id,
    [msg(From, Id1, response(Resp))|UserIn],
@@
    agent(Id?, UserIn?, NetIn1?, Outs1?,
        Holdings?, NextSerial?).
%% Cold-call from network
agent(Id, UserIn,
    [msg(Id1, intro(From, Resp?))|NetIn],
@@
    agent(Id?, UserIn?, NetIn?, Outs1?,
        Holdings?, NextSerial?).
%% Trade command
agent(Id, [msg(’_user’, Id1,
@@
        Holdings?, UserIn?, NetIn?, Outs?,
        NextSerial?).
%% Trade complete (injected — accept)
agent(Id,
    [trade_complete(From, TheirBonds)|UserIn],
@@
    ground(Id?), ground(From?),
    ground(TheirBonds?) |
@@
    agent(Id?, UserIn?, NetIn?, Outs1?,
        NewHoldings?, NextSerial?).
%% Trade returned (injected — decline)
agent(Id,
    [trade_returned_bonds(From, OurBonds)|UserIn],
@@
    ground(Id?), ground(From?),
    ground(OurBonds?) |
@@
    agent(Id?, UserIn?, NetIn?, Outs1?,
        NewHoldings?, NextSerial?).
%% Trade returned with menu
agent(Id, [trade_returned_bonds_menu(From,
    OurBonds, Menu)|UserIn],
@@
    agent(Id?, UserIn?, NetIn?, Outs1?,
        NewHoldings?, NextSerial?).
%% Incoming trade_propose
agent(Id, UserIn, [msg(From, Id1,
@@
    classify_trade(Id?, WantSpec?,
        OfferedBonds?, TradeClass),
    trade_dispatch(TradeClass?, Id?, From?,
        WantSpec?, OfferedBonds?, TradeResp,
        Holdings?, UserIn?, NetIn?, Outs?,
        NextSerial?).
%% Accept trade
agent(Id, [msg(’_user’, Id1,
    accept_trade(From,
@@
    handle_trade_fill(Status?, Id?, From?,
        TradeResp, OfferedBonds?, Selected?,
        Remaining?, UserIn?, NetIn?, Outs?,
        NextSerial?).
%% Reject trade
agent(Id, [msg(’_user’, Id1,
    reject_trade(From,
        trade_pending(TradeResp?, _,
            OfferedBonds)))|UserIn],
@@
    Id? =?= Id1?, ground(From?),
    ground(OfferedBonds?) |
@@
    agent(Id?, UserIn?, NetIn?, Outs?,
        Holdings?, NextSerial?).
%% Deposit escrow
agent(Id, [msg(’_user’, Id1,
@@
        ReleaseTime?, Holdings?, UserIn?, NetIn?,
        Outs?, NextSerial?).
%% Escrow expired (injected)
@@
        msg(agent, ’_user’,
            escrow_expired(Target?)),
@@
    agent(Id?, UserIn?, NetIn?, Outs1?,
        Holdings?, NextSerial?).
%% Escrow returned (injected)
@@
    ground(Id?), ground(Target?),
    ground(OurBonds?) |
@@
        msg(agent, ’_user’,
            escrow_returned(Target?)),
@@
    agent(Id?, UserIn?, NetIn?, Outs1?,
        NewHoldings?, NextSerial?).
%% Cancel escrow
agent(Id, [msg(’_user’, Id1,
    cancel_escrow(Target,
@@
    agent(Id?, UserIn?, NetIn?, Outs?,
        Holdings?, NextSerial?).
%% Incoming escrow offer
agent(Id, UserIn, [msg(From, Id1,
@@
    Id? =?= Id1?, ground(From?),
    ground(ReleaseTime?) |
@@
    agent(Id?, UserIn1?, NetIn?, Outs1?,
        Holdings?, NextSerial?).
%% Escrow released (injected)
agent(Id,
    [escrow_ben_released(From, Bonds)|UserIn],
@@
        msg(agent, ’_user’,
            escrow_released(From?)),
@@
    agent(Id?, UserIn?, NetIn?, Outs1?,
        NewHoldings?, NextSerial?).
%% Escrow cancelled (injected)
@@
        msg(agent, ’_user’,
            escrow_cancelled(From?)),
@@
    agent(Id?, UserIn?, NetIn?, Outs1?,
        Holdings?, NextSerial?).
%% Termination
agent(_, [], _, Outs, _, _) :-
    close_outputs(Outs?).